OpenStudy (anonymous):

the are of a circle with radius R is given by formula a=pieR^2. if circle has area 36 pie sq inches what is the radiusin inches

OpenStudy (phi):

They are saying \( Area= \pi r^2 \) and \( Area= 36 \pi \) square inches so you can write down the equation \( \pi r^2 = 36 \pi \) sq. in can you solve for r^2 and then r?

OpenStudy (anonymous):

my problem is i dont ahve a calculator with the pie symbol tho

OpenStudy (anonymous):

would it be like 6 pie?

OpenStudy (phi):

If you start with \[ \pi r^2 = 36 \pi \] what happens if you divide both sides by \( \pi \)?

OpenStudy (anonymous):

r^2=36 r=6?

OpenStudy (phi):

yes

OpenStudy (phi):

\[ \frac{\cancel{\pi} r^2}{\cancel{\pi}} =\frac{\cancel{\pi} 36}{\cancel{\pi}} \] \[ r= \pm 6 \] but r is always positive for a circle r=6

OpenStudy (anonymous):

thank u